OpenStudy (anonymous):

Harry is trying to solve the equation y = 2x2 − x − 6 using the quadratic formula. He has made an error in one of the steps below. Find the step where Harry went wrong. please solve i need help Harry is trying to solve the equation y = 2x2 − x − 6 using the quadratic formula. He has made an error in one of the steps below. Find the step where Harry went wrong. x= 1 plus or minus the square root of forty sevenall over two times

jhonyy9 (jhonyy9):

2x^2 -x -6=0 D=1-4*2*(-6)=1+48=49 x_1,2=(1+/- sqrt(49) )/4 = (1+/-7)/4 =-6/4 =-3/2 and 8/4 = 2

jhonyy9 (jhonyy9):

so the discriminant was calculated wrong not is 47 the discriminant is 49

jhonyy9 (jhonyy9):

ok ?